# Preasymptotic error estimates of higher-order EEM for the time-harmonic Maxwell equations with large wave number

### v1

@@ Abstract.
This work was partially supported by the NSF of China under grants 12171238, 12261160361, and 11525103.
@@ 1. Introduction
$\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}\bm{E}-\kappa^{2}\bm{E}=\bm{f}$ in $\quad\text{in }\Omega,$ (1.1a)
@@
A major challenge when solving high-frequency wave scattering problems is that FEM of fixed order may suffer from the “pollution effect”, where accuracy deteriorates as the wave number increases [7, 30, 31]. This issue is particularly significant in both theoretical and practical applications, necessitating the derivation of error estimates that explicitly account for the impact of $\kappa$. In what follows, by “asymptotic error estimate”, we mean an wave-number-explicit error estimate without pollution error (i.e., quasi-optimal); in contrast, by “preasymptotic error estimate”, we mean one with nonnegligible pollution effect. The understanding of FEM for the high-frequency Helmholtz problem $\Delta u+\kappa^{2}u=f$ has seen significant advancements over the past few decades through various works [7, 30, 31, 40, 41, 57, 21, 55, 56, 33, 24, 16, 34]. For example, it has been shown in [21, 24] that when applying FEM to the Helmholtz equation with an impedance boundary condition, the following preasymptotic error bound holds if $\kappa^{2p+1}h^{2p}$ is sufficiently small:
@@
where $\interleave\cdot\interleave_{h}=\big{(}\left\|\nabla\cdot\right\|^{2}+\kappa^{2}\|\cdot\|^{2}\big{)}^{\frac{1}{2}}$, $V_{h}$ is the $p$-th order Lagrange finite element space and $u_{h}$ is the discrete solution. The above error bound (1.2) shows that the actual error of FEM is the sum of the best approximation error plus a pollution term which grows rapidly when $\kappa\rightarrow+\infty$. On one hand, preasymptotic error estimates like (1.2) provide valuable insights into how the wave number $\kappa$ influence the reliability and accuracy of FEM as $\kappa$ increases. And one can deduce that $\kappa^{p+1}h^{p}\lesssim 1$ ensures the quasi-optimality of $p$-th order FEM. On the other hand, the above mesh condition is practical, as achieving a discrete solution with reasonable accuracy requires the pollution term $\mathcal{O}(\kappa^{2p+1}h^{2p})$ to be small. Both works [21] and [24] are based on the modified duality argument framework (cf. [55, 56, 57]) to obtain preasymptotic error estimates (1.2). However, they employ different techniques to obtain results under the weakest currently known mesh condition, $\kappa^{2p+1}h^{2p}\lesssim 1$. The work in [21] uses the negative-norm estimates technique, while the authors of [24] adapt the traditional elliptic-projection for Helmholtz-type problems (cf. [55, 57]) by constructing a novel elliptic operator, which can be obtained by adding an appropriate smoothing operator to the original partial differential operator.
@@
Now, a natural and important question arises: how does the accuracy deteriorate as $\kappa\rightarrow+\infty$ when the EEM is applied to the time-harmonic Maxwell equations? Or more precisely, does a similar preasymptotic error bound (1.2) still hold for the Maxwell cases? Compared to high-frequency Helmholtz-type problems, analyses for indefinite Maxwell problems are more intricate because of the nontrivial kernel space of the $\operatorname{\bm{\mathrm{curl}}}$ operator. Up to now, the $\kappa$-implicit convergence analysis (see, e.g., [46, 25, 47]) and asymptotic error estimates of EEM for (1.1a) are well-established (see, e.g., [42, 43, 13]) and existing dispersion analysis on structured meshes (see, e.g., [2, 3, 4]) suggest that the EEM applied to the Maxwell problems has similar dispersive behavior as that of the FEM. Among these works, a milestone is the asymptotic error estimates proposed in [42, 43] for the time-harmonic Maxwell equations with transparent and impedance boundary conditions, respectively. They stated that the discrete solution of $hp$-EEM is pollution-free (i.e., quasi-optimal) if $p\geq\max\{1,C\ln\kappa\}$ for any $C>0$ and $\kappa h/p$ is sufficiently small. For the fixed order EEM (note that $p=0$ corresponds to the lowest edge elements in [42, 43]), it requires that $\kappa^{p+7}h^{p}$ is sufficiently small (see [42, Remark 4.19]) for the case of transparent boundary condition or $\kappa^{p+1}h^{p}$ is sufficiently small (see [43, Lemma 9.5]) for the case of impedance boundary condition. However, rigorous preasymptotic error estimates for EEM applied to time-harmonic Maxwell equations remain scarce in the literature. Addressing the above question forms the primary motivation of this paper.
@@
under the mesh condition $\kappa^{2p+2}h^{2p}C_{\rm sol}\leq C_{2}$ for some positive constants $C_{1}$ and $C_{2}$. Here, $C_{\rm sol}$ is the ${\bm{L}}^{2}(\Omega)\to{\bm{L}}^{2}(\Omega)$ norm of the solution operator $\bm{f}\mapsto\bm{E}$, and $\bm{V}_{h}$ represents the first- or second-type edge element space of order $p$. This work extends the methodology proposed in [24] to the time-harmonic Maxwell problems with PEC boundary conditions. A key insight from [14] is that the analogous elliptic operator in[24] for Maxwell problems can be constructed by adding a smoothing operator for functions lying in the complement of the kernel space of the $\operatorname{\bm{\mathrm{curl}}}$ operator. However, the explicit dependence of the stability constant $C_{\rm sol}$ on $\kappa$ was not considered in [14].
In this paper, a sequel of our previous work [38], we study the higher-order EEM for the time-harmonic Maxwell equations (1.1) and prove the following preasymptotic error estimate under the condition that $\kappa^{2p+1}h^{2p}$ is sufficiently small:
@@
where $\bm{V}_{h}$ is the second-type edge element space of order $p$ and ${\bm{E}}_{h}\in\bm{V}_{h}$ is the EEM solution. Note that the Assumption 2.3 (ii) in [14] relies on the fact that functions in ${\bm{H}}_{0}(\operatorname{\bm{\mathrm{curl}}};\Omega)$ have vanishing tangential trace. However, similar regularity results, such as those in [14, Lemma 11.2, Lemma 11.5] do not hold for the system (1.1). Thus the methodology presented in [14] cannot be applied to the analysis of problem (1.1). To address this limitation, we drew inspiration from the regularization operators employed in [24, 34, 14] and developed new preasymptotic error estimates for higher-order EEM, building on our previous work [38]. Specifically, we constructed a regularization operator tailored for (1.1a) with impedance boundary conditions. This operator effectively bridges the gap in regularity requirements, enabling us to refine the analysis of higher-order EEM to problem (1.1). Rather than constructing the regularization operator abstractly through functional calculus, as done in [24] and [14], we define the smoothing operator explicitly by truncating eigenfunction expansions. This straightforward and explicit construction is crucial for preasymptotic error estimation when higher-order EEM is applied to (1.1), allowing readers to focus on the key differences between the methodology used in [38] and this work. Moreover, our preasymptotic error bounds do not contains any $\kappa$-dependent constants. The key contributions of this work are summarized as follows:
@@
  • Wave-number-explicit ${\bm{H}}^{m}$ regularity results for (1.1) are derived under the assumption of a $C^{m}$-domain. These results improve upon those in [43, 36].
@@
  • The preasymptotic error estimate (1.5) in the energy norm is proved under the mesh condition that $\kappa^{2p+1}h^{2p}$ is sufficiently small. Using the second-type edge element method, this result generalizes the well-known preasymptotic error bound (1.2) to the more complicated time-harmonic Maxwell equations with impedance boundary condition.
@@
  • Preasymptotic error estimate in the ${\bm{L}}^{2}$ norm is also obtained under the same mesh condition.
@@
The remainder of this paper is organized as follows. In Section 2, we initially present the variational formulation of (1.1) and formulate the EEM. Subsequently, we furnish preliminary results related to stability, regularity, and approximation estimates. Section 3 introduces a particular auxiliary problem by utilizing a custom-designed smoothing operator. In Section 4, we first present some dual approximation results and the elliptic projection error estimates. Thereafter, we derive the preasymptotic error estimates for the EEM. Finally, in Section 5, we conduct numerical experiments to illustrate our theoretical findings.
@@
Throughout this paper, we use notations $A\lesssim B$ and $A\gtrsim B$ for the inequalities $A\leq CB$ and $A\geq CB$ respectively, where $C$ is a positive number independent of the mesh size and wave number $\kappa$, but the value of which may vary in different instances. The shorthand notation $A\simeq B$ represents $A\leq CB$ and $B\leq CA$ hold simultaneously. We suppose $\lambda\simeq 1$ and $\kappa\gg 1$ since we consider high frequency problem. Let =$\mathbb{N}=\{0,1,2,\cdots\}$ and $\mathbb{N}_{1}=\mathbb{N}\setminus\{0\}$. In this paper, we assume that the bounded open set $\Omega\in\mathbb{R}^{3}$ is strictly star-shaped with respect to a point ${\bm{x}}_{0}\in\Omega$, that is, there exists $C>0$ such that $({\bm{x}}-{\bm{x}}_{0})\cdot\bm{\nu}\geq C\;\;\forall\bm{x}\in\Gamma$. We remark that in this setting, $\Gamma=\partial\Omega$ is connected. Without loss of generality, we assume that $\operatorname{{diam}}(\Omega)\simeq 1$ since the case $\operatorname{{diam}}(\Omega)\gtrsim 1$ is equivalent to a large wave number problem through a suitable coordinate scaling.
@@ 2.1. Spaces and norms
To state the variational formulation of (1.1), we first introduce some notations. For a bounded domain $D\subset\mathbb{R}^{3}$ with a Lipschitz boundary $\partial D$, and any real number $s\geq 0$, we shall use the standard Sobolev space $H^{s}(D)$ with norm $\|\cdot\|_{s,D}$, and refer to [1, 8, 46] for their definitions. On the boundary of $D$, the space $H^{s}(\partial D)$ is well-defined if $D$ is $C^{m}$ for some $m\in\mathbb{N}_{1}$ and $s<m$ (cf., [5]). For vector-valued functions, these spaces are indicated by boldface symbols, such as ${\bm{H}}^{s}(D)$ and ${\bm{H}}^{s}(\partial D)$. When $D=\Omega$, we drop the index $s$ if $s=0$ and use the shorthands $\|\cdot\|:=\|\cdot\|_{0,\Omega},\|\cdot\|_{\Gamma}:=\|\cdot\|_{0,\Gamma}$ for simplicity. We denote by $(\cdot,\cdot)_{D}$ and $\langle\cdot,\cdot\rangle_{\Sigma}$ for $\Sigma\subset\partial D$ the ${\bm{L}}^{2}$ inner-product on $D$ and $\partial D$, respectively. Define $(\cdot,\cdot):=(\cdot,\cdot)_{\Omega}$ and $\langle\cdot,\cdot\rangle:=\langle\cdot,\cdot\rangle_{\Gamma}$. We introduce the following $\kappa$-weighted norms $\left\|\cdot\right\|_{m}$ on ${\bm{H}}^{m}(\Omega)$ for $m\in\mathbb{N}$ and $\left\|\cdot\right\|_{m-\frac{1}{2},\Gamma}$ on ${\bm{H}}^{m-\frac{1}{2}}(\Gamma)$ for $m\in\mathbb{N}_{1}$, respectively:
@@
$H_{0}^{1}(\Omega)$ :=$:=\{v\in H^{1}(\Omega):v|_{\Gamma}=0\},$
$\bm{H}(\operatorname{\bm{\mathrm{curl}}};\Omega)$ :=$:=\{\bm{v}\in{\bm{L}}^{2}(\Omega):\operatorname{\bm{\mathrm{curl}}}\bm{v}\in{\bm{L}}^{2}(\Omega)\},$
$\bm{H}^{s}(\operatorname{\bm{\mathrm{curl}}};\Omega)$ :=$:=\{\bm{v}\in{\bm{H}}^{s}(\Omega):\operatorname{\bm{\mathrm{curl}}}\bm{v}\in{\bm{H}}^{s}(\Omega)\},$
$\bm{H}_{0}(\operatorname{\bm{\mathrm{curl}}};\Omega)$ :=$:=\{\bm{v}\in\bm{H}(\operatorname{\bm{\mathrm{curl}}};\Omega):\bm{v}\times\bm{\nu}|_{\Gamma}=\bm{0}\},$
$\bm{H}(\operatorname{{\mathrm{div}}};\Omega)$ :=$:=\{\bm{v}\in{\bm{L}}^{2}(\Omega):\operatorname{{\mathrm{div}}}\bm{v}\in L^{2}(\Omega)\},$
$\bm{H}_{0}(\operatorname{{\mathrm{div}}};\Omega)$ :=$:=\{\bm{v}\in\bm{H}(\operatorname{{\mathrm{div}}};\Omega):\bm{v}\cdot\bm{\nu}|_{\Gamma}=0\},$
$\bm{H}(\operatorname{{\mathrm{div}}}^{0};\Omega)$ :=$:=\{\bm{v}\in{\bm{L}}^{2}(\Omega):\operatorname{{\mathrm{div}}}\bm{v}=0\},$
$\bm{H}^{s}(\operatorname{{\mathrm{div}}}^{0};\Omega)$ :=$:=\{\bm{v}\in{\bm{H}}^{s}(\Omega):\operatorname{{\mathrm{div}}}\bm{v}=0\},$
@@
$\bm{H}^{s}(\operatorname{{\mathrm{div}}}_{\Gamma};\Gamma)$ :=$:=\{\bm{v}\in{\bm{H}}^{s}(\Gamma):\operatorname{{\mathrm{div}}}_{\Gamma}\bm{v}\in{\bm{H}}^{s}(\Gamma)\},$
${\bm{L}}^{2}_{t}(\Gamma)$ :=$:=\{\bm{v}\in{\bm{L}}^{2}(\Gamma):\bm{v}\cdot\bm{\nu}|_{\Gamma}=0\},$

### v2

@@ Abstract.
This work was partially supported by National Key R&D Program of China 2024YFA1012600, and by the NSF of China under grants 12171238 and 12261160361.
@@ 1. Introduction
$\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}\bm{E}-\kappa^{2}\bm{E}=\bm{f}$ $\quad\text{in }\Omega,$ (1.1a)
@@
A major challenge when solving high-frequency wave scattering problems is that FEM of fixed order may suffer from the “pollution effect”, where accuracy deteriorates as the wave number increases [7, 30, 31]. This issue is particularly significant in both theoretical and practical applications, necessitating the derivation of error estimates that explicitly account for the impact of $\kappa$. In what follows, by “asymptotic error estimate”, we mean a wave-number-explicit error estimate without pollution error (i.e., quasi-optimal); in contrast, by “preasymptotic error estimate”, we mean one with nonnegligible pollution effect. The understanding of FEM for the high-frequency Helmholtz problem $\Delta u+\kappa^{2}u=f$ has seen significant advancements over the past few decades through various works [7, 30, 31, 40, 41, 58, 21, 56, 57, 33, 24, 16, 34]. For example, it has been shown in [21, 24] that when applying FEM to the Helmholtz equation with an impedance boundary condition, the following preasymptotic error bound holds if $\kappa^{2p+1}h^{2p}$ is sufficiently small:
@@
where $\interleave\cdot\interleave_{h}=\big(\left\|\nabla\cdot\right\|^{2}+\kappa^{2}\|\cdot\|^{2}\big)^{\frac{1}{2}}$, $V_{h}$ is the $p$-th order Lagrange finite element space and $u_{h}$ is the discrete solution. The above error bound (1.2) shows that the actual error of FEM is the sum of the best approximation error plus a pollution term which grows rapidly when $\kappa\rightarrow+\infty$. On one hand, preasymptotic error estimates like (1.2) provide valuable insights into how the wave number $\kappa$ influence the reliability and accuracy of FEM as $\kappa$ increases. And one can deduce that the quasi-optimality of $p$-th order FEM is ensured by the condition that $\kappa^{p+1}h^{p}$ is sufficiently small. On the other hand, the above mesh condition is practical, as achieving a discrete solution with reasonable accuracy requires the pollution term $\mathcal{O}(\kappa^{2p+1}h^{2p})$ to be small. Both works [21] and [24] are based on the idea of modified duality argument [58] to obtain preasymptotic error estimates (1.2). However, they employ different techniques to obtain results under the weakest currently known mesh condition that $\kappa^{2p+1}h^{2p}$ is sufficiently small. The work in [21] uses the traditional elliptic-projection combined with the negative-norm estimates technique, while the authors of [24] adapt the traditional elliptic-projection by constructing a novel elliptic operator, which can be obtained by adding an appropriate smoothing operator to the original partial differential operator.
@@
Now, a natural and important question arises: how does the accuracy deteriorate as $\kappa\rightarrow+\infty$ when the EEM is applied to the time-harmonic Maxwell equations? Or more precisely, does a similar preasymptotic error bound (1.2) still hold for the Maxwell cases? Compared to high-frequency Helmholtz-type problems, analyses for indefinite Maxwell problems are more intricate because of the nontrivial kernel space of the $\operatorname{\bm{\mathrm{curl}}}$ operator. Up to now, the $\kappa$-implicit convergence analysis (see, e.g., [47, 25, 48]) and asymptotic error estimates of EEM for (1.1a) are well-established (see, e.g., [42, 43, 13]) and existing dispersion analysis on structured meshes (see, e.g., [2, 3, 4]) suggest that the EEM applied to the Maxwell problems has similar dispersive behavior as that of the FEM for the Helmholtz problem. Among these works, a milestone is the asymptotic error estimates proposed in [42, 43] for the time-harmonic Maxwell equations with transparent and impedance boundary conditions, respectively. They proved that the discrete solution of $hp$-EEM is pollution-free (i.e., quasi-optimal) under the condition that $p\geq C\ln\kappa$ for some $C>0$ and $\kappa h/p$ is sufficiently small. In contrast, for the fixed order EEM, for instance, in the case of impedance boundary condition, it requires that $\kappa^{p}h^{p-1}$ is sufficiently small (see [43, Lemma 9.5], note that $p=2$ here corresponds to $p=1$ in that reference, which refers to the quadratic edge element of the first type). However, rigorous preasymptotic error estimates for EEM applied to time-harmonic Maxwell equations remain incomplete in the literature. Addressing the above question forms the primary motivation of this paper.
@@
under the mesh condition $\kappa^{2p+2}h^{2p}C_{\rm sol}\leq C_{2}$ for some positive constants $C_{1}$ and $C_{2}$. Here, $C_{\rm sol}$ is the ${\bm{L}}^{2}(\Omega)\to{\bm{L}}^{2}(\Omega)$ norm of the solution operator $\bm{f}\mapsto\bm{E}$, and $\bm{V}_{h}$ represents the first- or second-type edge element space of order $p$. This work extends the methodology proposed in [24] to the time-harmonic Maxwell problems with PEC boundary conditions. A key insight from [14] is that the analogous elliptic operator in[24] for Maxwell problems can be constructed by adding a smoothing operator for functions lying in the complement of the kernel space of the $\operatorname{\bm{\mathrm{curl}}}$ operator.
In this paper, a sequel of our previous work [38], we study the higher-order EEM for the time-harmonic Maxwell problem (1.1) with impedance boundary condition and prove the following preasymptotic error estimate under the condition that $\kappa^{2p+1}h^{2p}$ is sufficiently small:
@@
where $\bm{V}_{h}$ is the second-type edge element space of order $p$ and ${\bm{E}}_{h}\in\bm{V}_{h}$ is the EEM solution. Note that the methodology presented in [14] for the Maxwell problem with PEC boundary condition cannot be directly applied to the analysis of problem (1.1). For example, Assumption 2.3 (ii) in [14] relies on the fact that functions in ${\bm{H}}_{0}(\operatorname{\bm{\mathrm{curl}}};\Omega)$ have vanishing tangential trace, and the regularity results established in [14, Lemma 11.2, Lemma 11.5] do not hold for system (1.1). To overcome this limitation and develop new preasymptotic error estimates for higher-order EEM for the problem (1.1), we first drew inspiration from the auxiliary elliptic operators used in [24, 34, 14], and constructed a new Maxwell elliptic operator tailored for the Maxwell problem (1.1) with impedance boundary conditions, using a smoothing operator defined simply by truncating eigenfunction expansions. We then provide a new and simpler proof of the regularity decomposition result from [43, Theorem 7.3] but under a weaker smoothness assumption on the domain $\Omega$. Based on this decomposition result and using the trick of “modified duality argument” [58, 38, 14, 34], we may prove that the EEM satisfies the same stability estimates (see Theorem 4.2) as the continuous problem (1.1) (see (2.17)), under the mesh condition that $\kappa^{2p+1}h^{2p}$ is sufficiently small. Finally, we derive the preasymptotic error estimates for the EEM in both energy and ${\bm{L}}^{2}$ norms (see (1.5) and Theorem 4.3). Moreover, our preasymptotic error bounds do not contain any $\kappa$-dependent constants. The key contributions of this work are summarized as follows:
@@
  • Wave-number-explicit ${\bm{H}}^{m}$ regularity and regularity by decomposition results (see Theorems 2.1 and 4.1 for the problem (1.1)) are derived under the assumption of a $C^{m}$-domain. These results improve upon those in [43, 36].
@@
  • The preasymptotic error estimate (1.5) in the energy norm is proved under the mesh condition that $\kappa^{2p+1}h^{2p}$ is sufficiently small. Using the second-type edge element method, this result generalizes the well-known preasymptotic error bound (1.2) for the Helmholtz problem to the more challenging time-harmonic Maxwell equations with impedance boundary condition.
@@
  • Preasymptotic error estimate in the ${\bm{L}}^{2}$ norm is also obtained under the same mesh condition. This is the first such result for higher-order EEM of the second type applied to time harmonic Maxwell problems.
@@
The remainder of this paper is organized as follows. In Section 2, we initially present the variational formulation of (1.1) and formulate the EEM. Subsequently, we furnish preliminary results related to stability, regularity, and approximation estimates. Section 3 introduces the auxiliary Maxwell elliptic problem and proves error estimates for the corresponding elliptic projections. In Section 4, we prove the decomposition result and derive the preasymptotic stability and error estimates for the EEM. Finally, in Section 5, we conduct numerical experiments to illustrate our theoretical findings.
@@
Throughout this paper, we use notations $A\lesssim B$ and $A\gtrsim B$ for the inequalities $A\leq CB$ and $A\geq CB$ respectively, where $C$ is a positive number independent of the mesh size and wave number $\kappa$, but the value of which may vary in different instances. The shorthand notation $A\simeq B$ represents $A\leq CB$ and $B\leq CA$ hold simultaneously. We suppose $\lambda\simeq 1$ and $\kappa\gg 1$ since we consider high frequency problem. Let $\mathbb{N}=\{0,1,2,\cdots\}$ and $\mathbb{N}_{1}=\mathbb{N}\setminus\{0\}$. In this paper, we assume that the bounded open set $\Omega\in\mathbb{R}^{3}$ is strictly star-shaped with respect to a point ${\bm{x}}_{0}\in\Omega$, that is, there exists $C>0$ such that $({\bm{x}}-{\bm{x}}_{0})\cdot\bm{\nu}\geq C\;\;\forall\bm{x}\in\Gamma$. We remark that in this setting, $\Gamma=\partial\Omega$ is connected. Without loss of generality, we assume that $\operatorname{{diam}}(\Omega)\simeq 1$ since the other case of $\operatorname{{diam}}(\Omega)\not\simeq 1$ can be transformed to this case through a suitable coordinate scaling.
@@ 2.1. Spaces and norms
To state the variational formulation of (1.1), we first introduce some notations. For a bounded domain $D\subset\mathbb{R}^{3}$ with a Lipschitz boundary $\partial D$, and any real number $s\geq 0$, we shall use the standard Sobolev spaces $L^{\infty}(D)$ and $H^{s}(D)$, their norms $\|\cdot\|_{L^{\infty}(D)}$ and $\|\cdot\|_{s,D}$, and refer to [1, 8, 47] for their definitions. On the boundary of $D$, the space $H^{s}(\partial D)$ is well-defined if $D$ is $C^{m}$ for some $m\in\mathbb{N}_{1}$ and $s<m$ (cf., [5]). For vector-valued functions, these spaces are indicated by boldface symbols, such as ${\bm{H}}^{s}(D)$ and ${\bm{H}}^{s}(\partial D)$. When $D=\Omega$, we omit the subscript $D$ and the index $s$ if $s=0$ and use the shorthands $\|\cdot\|:=\|\cdot\|_{0,\Omega},\|\cdot\|_{\Gamma}:=\|\cdot\|_{0,\Gamma}$ for simplicity. We denote by $(\cdot,\cdot)_{D}$ and $\langle\cdot,\cdot\rangle_{\Sigma}$ for $\Sigma\subset\partial D$ the ${\bm{L}}^{2}$ inner-product on $D$ and $\partial\Sigma$, respectively. For simplicity, write $(\cdot,\cdot):=(\cdot,\cdot)_{\Omega}$ and $\langle\cdot,\cdot\rangle:=\langle\cdot,\cdot\rangle_{\Gamma}$. We introduce the following $\kappa$-weighted norms $\left\|\cdot\right\|_{m,\kappa}$ on ${\bm{H}}^{m}(\Omega)$ for $m\in\mathbb{N}$ and $\left\|\cdot\right\|_{m-\frac{1}{2},\kappa,\Gamma}$ on ${\bm{H}}^{m-\frac{1}{2}}(\Gamma)$ for $m\in\mathbb{N}_{1}$, respectively:
@@
$H_{0}^{1}(\Omega)$ $:=\{v\in H^{1}(\Omega):v|_{\Gamma}=0\},$
$\bm{H}(\operatorname{\bm{\mathrm{curl}}};\Omega)$ $:=\{\bm{v}\in{\bm{L}}^{2}(\Omega):\operatorname{\bm{\mathrm{curl}}}\bm{v}\in{\bm{L}}^{2}(\Omega)\},$
$\bm{H}^{s}(\operatorname{\bm{\mathrm{curl}}};\Omega)$ $:=\{\bm{v}\in{\bm{H}}^{s}(\Omega):\operatorname{\bm{\mathrm{curl}}}\bm{v}\in{\bm{H}}^{s}(\Omega)\},$
$\bm{H}_{0}(\operatorname{\bm{\mathrm{curl}}};\Omega)$ $:=\{\bm{v}\in\bm{H}(\operatorname{\bm{\mathrm{curl}}};\Omega):\bm{v}\times\bm{\nu}|_{\Gamma}=\bm{0}\},$
$\bm{H}(\operatorname{{\mathrm{div}}};\Omega)$ $:=\{\bm{v}\in{\bm{L}}^{2}(\Omega):\operatorname{{\mathrm{div}}}\bm{v}\in L^{2}(\Omega)\},$
$\bm{H}_{0}(\operatorname{{\mathrm{div}}};\Omega)$ $:=\{\bm{v}\in\bm{H}(\operatorname{{\mathrm{div}}};\Omega):\bm{v}\cdot\bm{\nu}|_{\Gamma}=0\},$
$\bm{H}(\operatorname{{\mathrm{div}}}^{0};\Omega)$ $:=\{\bm{v}\in{\bm{L}}^{2}(\Omega):\operatorname{{\mathrm{div}}}\bm{v}=0\},$
$\bm{H}^{s}(\operatorname{{\mathrm{div}}}^{0};\Omega)$ $:=\{\bm{v}\in{\bm{H}}^{s}(\Omega):\operatorname{{\mathrm{div}}}\bm{v}=0\},$
@@
$\bm{H}^{s}(\operatorname{{\mathrm{div}}}_{\Gamma};\Gamma)$ $:=\{\bm{v}\in{\bm{H}}^{s}(\Gamma):\operatorname{{\mathrm{div}}}_{\Gamma}\bm{v}\in H^{s}(\Gamma)\},$
${\bm{L}}^{2}_{t}(\Gamma)$ $:=\{\bm{v}\in{\bm{L}}^{2}(\Gamma):\bm{v}\cdot\bm{\nu}|_{\Gamma}=0\},$